# Full-body WPT: wireless powering with meandered e-textiles

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Interaction devicesccs: Human-centered computing Ubiquitous and mobile computingccs: Hardware Wireless devices
@@ 5. Power transfer efficiency
Figure 5. Illustration of Full-body Internet of Textiles using textile-based near-field power and data networks. Our textile-based network enables to wirelessly send power and data to multiple on-body devices around the body.
@@ 6. Application examples
7. Discussion
Meander coils can be fabricated using a range of conductive materials, each with trade-offs between efficiency, weight, and ease of integration. Liquid metal, for example, offers low resistance and high power transfer efficiency. However, its use requires thick tubing, resulting in substantial added weight—often exceeding 100 grams—which limits its practicality for daily wearable applications. In contrast, conductive yarns can be machine-knitted directly into garments, providing excellent flexibility and low weight. Yet, their higher electrical resistance leads to significant power loss, reducing overall efficiency to around 5%. To address this trade-off, we adopt copper-coated fabric as the conductor for the meander coil. This material combines the benefits of low resistance—achieving efficiency comparable to liquid metal—with lightweight, textile-compatible form factors.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Companion of the the 2025 ACM International Joint Conference on Pervasive and Ubiquitous Computing; October 12–16, 2025; Espoo, Finlandbooktitle: Companion of the the 2025 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp Companion ’25), October 12–16, 2025, Espoo, Finlanddoi: 10.1145/3714394.3750714isbn: 979-8-4007-1477-1/2025/10ccs: Human-centered computing Interaction devicesccs: Human-centered computing Ubiquitous and mobile computingccs: Hardware Wireless devices
@@ 5. Power transfer efficiency
Figure 5. Illustration of Full-body Internet of Textiles using textile-based near-field power and data networks.